OpenStudy (anonymous):

One number is five more than three times another. If their difference is 35, what is the larger number? 15 50 65

Parth (parthkohli):

One number is 5 more than three times another ----> 5 + 3x Another ----> x Equation would look like: \( \color{Black}{\Rightarrow (5 + 3x) - x = 35 }\)

Parth (parthkohli):

Remove the brackets. They're of no use here.

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

do you still need help?

OpenStudy (anonymous):

how do i plug thaat into my calculator to get the answer

Parth (parthkohli):

Lol you need a calculator for that?

OpenStudy (anonymous):

i get 7 otherwise

Parth (parthkohli):

7? How?

OpenStudy (lgbasallote):

you divided by 5? no no....just combine 5 + 3x - x = 35

Parth (parthkohli):

What is 3x - x?

OpenStudy (lgbasallote):

3x and -x are common terms so combine them...what do you get?

OpenStudy (anonymous):

2x

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5 + 2x = 35 }\) What's the next step now?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

from both sides

OpenStudy (lgbasallote):

good

OpenStudy (lgbasallote):

what does that give?

OpenStudy (anonymous):

2x=30

Parth (parthkohli):

Yes: \( \color{Black}{\Rightarrow 2x = 30 - 5}\) \( \color{Black}{\Rightarrow 2x = 30 }\) What's next?

OpenStudy (anonymous):

divide by 2?

Parth (parthkohli):

Yes, and what does that give?

OpenStudy (lgbasallote):

GOOD!!!

OpenStudy (lgbasallote):

so x is....?

OpenStudy (anonymous):

so u get 5!!!

Parth (parthkohli):

Now plug in the value of x in 5 + 2x

OpenStudy (anonymous):

*15

OpenStudy (lgbasallote):

YOU GOT IT!!!!

Parth (parthkohli):

5 + 3x

Parth (parthkohli):

What is \( \color{Black}{\Rightarrow 5 + 3(15) }\)?

OpenStudy (lgbasallote):

^that's to find the larger number

OpenStudy (lgbasallote):

note that 15 is just the smaller number

Parth (parthkohli):

Yes, because I told in the first post that x is the smaller number and it's written that the larger is 5 more than 3 times the smaller one.

Parth (parthkohli):

Here, it'd be 5 more than 3 times of 15 which is 5 + 3(15)

OpenStudy (anonymous):

8(15)

OpenStudy (lgbasallote):

instead of that since y - x = 35 and x is 15 y - x = 35 y = 35 + 15 you can use that

Parth (parthkohli):

No not that one.

OpenStudy (lgbasallote):

you multiply 3 and 15 before adding 5

Parth (parthkohli):

\( \color{Black}{\Rightarrow (5) + [3(15)] }\)

OpenStudy (anonymous):

50

Parth (parthkohli):

Yes, you got it!

OpenStudy (lgbasallote):

in parth's equation i mean

Parth (parthkohli):

Correct :D

OpenStudy (lgbasallote):

50 is right :DDD

OpenStudy (anonymous):

thanks guys who should i give the medal to

Parth (parthkohli):

anyone, who deserves.

OpenStudy (anonymous):

@shivam_bhalla plzz help http://openstudy.com/study?signup#/updates/4fcb2338e4b0c6963ad4e149